OpenStudy (anonymous):

solve this problem::: 2(9-2)-3=

OpenStudy (angela210793):

2*7-3=14-3=11

OpenStudy (anonymous):

course im trying to check my answers

OpenStudy (anonymous):

the answer was 11 who put 13?? LOL